# The Nontrivial Vacuum Structure of an Extended $t\overline{t}$ BEH (Higgs) Bound state

### v1

@@ II Absence of Relative Time as a Gauge Invariance
We begin with the notion that the internal wave-function of two-body bound state, $\phi(r^{\mu})$, must not depend upon “relative time,” i.e., is independent of $r^{0}$ in the barycentric frame, or correspondingly $r^{\prime 0}$ in any frame. This can be seen in the free field limit by kinematics if we consider a pair of equal-mass particles of 4-momenta $p_{1}$ and $p_{2}$, $p_{1}^{2}=p_{2}^{2}=m^{2}$, and a bilocal wave-function, $H(x,y)\sim\exp(ip_{1}x+ip_{2}y)$. We pass to the total momentum $P=(p_{1}+p_{2})$ and relative momentum $Q=(p_{1}-p_{2})$, and the plane waves become $\exp(iPX+iQr)$, where we define “barycentric coordinates,” $x^{\mu}=X^{\mu}+r^{\mu}$, $y^{\mu}=X^{\mu}-r^{\mu}$. Note that $P_{\mu}Q^{\mu}=p_{1}^{2}-p_{2}^{2}=0$. Therefore, in the center-of-mass frame, in which $P=(P^{0},\vec{0})$ and $Q=(0,\vec{q})$, we see that $Q^{0}=0$. This implies there is no dependence in the bilocal state on $r^{0}$ through $\exp(iQ_{0}r^{0})=1$, and likewise no dependence upon a boosted relative time $r^{\prime 0}$ in any other frame. If the particles are constituents of a bound state, then the “relative time” must decouple from the dynamics.
@@
Given an arbitrary Lorentz invariant function, $\phi(r^{\mu})$, in any frame there will generally be dependence upon a relative time, $\sim\tau$. This is analogous to the gauge dependent components of a vector potential, and it is an artifact of using the bilocal field description. [^1] I am grateful to Bill Bardeen for some discussions that inspired this perspective. As described in the Introduction, the relative time, $\tau$, can be written in any given frame as $r^{\mu}=\omega^{\mu}\tau$, and implicitly requires the timelike, 4-vector, $\omega^{\mu}$, the “arrow of relative time.” Hence, eliminating dependence upon $\tau$, requires a Lorentz invariant constraint, such as $\omega_{\mu}\partial^{\mu}\phi(r)=0$.
In the symmetric phase of the standard model (SM), (or for any typical two-body bound state) the BEH boson contains such a vector, i.e., the 4-momentum $P^{\mu}$ carried by $H(X^{\mu})\sim\exp(iP_{\mu}X^{\mu})$. We can therefore bootstrap $\omega^{\mu}$ to $P^{\mu}$ through a constraint relation $\omega^{\mu}\propto P^{\mu}$. For example, we can do this semiclassically by introducing Lagrange multipliers into the action, such as:
$W=M_{0}^{4}\int d^{4}Xd^{4}r\;\lambda^{\prime}(H^{\dagger}D_{\mu}H\phi^{\dagger}\partial_{\mu}\phi)$ (9)
@@
Hence we demand $\delta W/\delta\lambda^{\prime}=0$, which imposes the kinematic constraint $P_{\mu}Q^{\mu}=0$. However, the deeper question remains: “what happens in the vacuum where $P_{\mu}=0$ and $\omega^{\mu}$ becomes unconstrained?”
@@
$\!\!\!\!\!\!\!\!\!\!S_{\phi}$ $\rightarrow$ $M_{0}^{4}\int\!\!d^{4}r\biggl(Z\partial_{\mu}\phi_{\omega}^{\dagger}(r)\partial^{\mu}\phi_{\omega}(r)+2g_{0}^{2}N_{c}D_{F}(2r^{\mu})\phi_{\omega}^{\dagger}(r)\phi_{\omega}(r)\biggr)$ (11)
@@
$M_{0}\int\!\!\biggl(-Z\frac{\partial^{2}\phi_{\omega}(r^{\mu})}{\partial r^{\mu}\partial r_{\mu}}+2g_{0}^{2}N_{c}D_{F}(2r^{\mu})\phi_{\omega}(r^{\mu})\biggr)\omega_{\nu}dr^{\nu}=ZM_{0}\int\!\!\mu^{2}\phi_{\omega}(r^{\mu})\omega_{\nu}dr^{\nu}=\mu^{2}\phi_{\omega}(r^{\mu})$ (12)
Note the presence of the overall line integral, $\int\omega_{\mu}dr^{\mu}$. This line integral remains in the equation of motion since $\phi_{\omega}(r^{\mu})$ has no dependence upon $r_{\mu}\propto\omega_{\mu}$, hence the variation is constrained, $\delta\phi_{\omega}\sim\delta^{3}(r^{\mu}_{\perp})$ where $\omega_{\mu}r^{\mu}_{\perp}=0$ and does not produce a longitudinal variation, $\delta(\omega_{\mu}r^{\mu}$). We define $Z$ by the line integral normalization chris1 ,
@@
It is important to realize that eq.(12) is not a conventional Klein-Gordon equation due to the line integral constraint.
@@
$\!\!\!\!\!\!\!\!\!\!S_{\phi}=M_{0}^{3}\!\!\int\!\!d^{3}r\;\biggl(-|\nabla_{\vec{r}}\phi(\vec{r})|^{2}+\int dr^{0}\;{2g_{0}^{2}N_{c}M_{0}}D_{F}(2r^{\mu})|\phi(\vec{r})\;|^{2}\biggr)$ (14)
@@
In the solution to the SKG equations with the eigenvalue, $\mu^{2}$, we see upon integrating by parts that $S_{\phi}\rightarrow\mu^{2}$, and the full bound state action then becomes:
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S=\!\!\int\!\!d^{4}X\;\biggl(|D_{H}H(X)|^{2}-\mu^{2}|H(X)|^{2}-\frac{\lambda}{2}(H^{\dagger}H)^{2}-g_{Y}\left([\overline{\psi}_{iL}(X)t_{R}(X)]_{f}H^{i}(X)+h.c.\right)\biggr)+S^{\prime}.$ (18)
$\mu^{2}$ is the physical mass of the bound state. The Yukawa coupling, $g_{Y}$ and quartic coupling $\lambda$ are defined in eqs.(79-84) in Appendix A. We emphasize that, while we have evaluated the action in the particular $\omega_{\mu}$ frame in eq.(14), this is just a calculation in a simplifying frame in the overall Lorentz invariant action eq.(12). The result eq.(18) holds in any frame.
@@
$\frac{g_{c}^{2}N_{c}}{8\pi^{2}}=1.06940.\qquad\makebox{c.f, the NJL critical value }\qquad\left.\frac{g_{0}^{2}N_{c}}{8\pi^{2}}\right|_{NJL}=1.00$ (19)
@@
$\overline{g}_{0}^{2}=g_{0}^{2}\left(1-\frac{g_{0}^{2}N_{c}}{8\pi^{2}}\right)^{-1}$ (20)
When $\overline{g}_{0}>g_{c}$ the eigenvalue is $\mu^{2}<0$. In such a solution the action eq.(18) for $H(X)$ then yields the “sombrero potential”:
@@ III.1 Formal Derivation
The action of eq.(11) is manifestly Lorentz invariant with dependence upon the 4-vector $\omega_{\mu}$. For the spherically symmetric ground state it must yield a Lorentz invariant expression in $\omega$. Hence, under a Lorentz transformation, $\omega^{\prime}_{\mu}=\Lambda_{\mu}^{\nu}\omega^{\nu}$ we can likewise perform $r^{\prime}_{\mu}=\Lambda_{\mu}^{\nu}r^{\nu}$, however, we are then free to change the integration variable back to its original form, $r^{\prime}\rightarrow r$ to obtain:
$\!\!\!\!\!\!\!\!\!\!S_{\phi}=M_{0}^{4}\int\!\!d^{4}r\biggl(Z\partial_{\mu}\phi_{\omega^{\prime}}^{\dagger}(r)\partial^{\mu}\phi_{\omega^{\prime}}(r)+2g_{0}^{2}N_{c}D_{F}(2r^{\mu})\phi_{\omega^{\prime}}^{\dagger}(r)\phi_{\omega^{\prime}}(r)\biggr)$ (22)
@@
$ZM_{0}^{4}\int\!\!d^{4}r\;\phi_{\omega}^{\dagger}(r^{\mu})F(r^{\mu})\phi_{\omega`}(r^{\mu})=ZM_{0}^{4}\int\!\!d^{4}r\;\delta^{4}(\omega-\omega^{\prime})\phi_{\omega}^{\dagger}(r^{\mu})F(r^{\mu})\phi_{\omega`}(r^{\mu})$ (23)
@@
$ZM_{0}^{4}\int\!\!d^{4}r\;\phi_{\omega}^{\dagger}(r^{\mu})\delta^{4}(r^{\mu})\phi_{\omega`}(r^{\mu})=ZM_{0}^{4}\phi_{\omega}^{\dagger}(0)\phi_{\omega`}(0)=ZM_{0}^{4}\int\!\!d^{4}r\;\delta^{4}(r^{\mu})\phi_{\omega}^{\dagger}(r)\phi_{\omega`}(r)$ (24)
This implies that we will have the BCS or BEC-like behavior for $\Phi(r)$. We derive the orthogonality in Appendix B.
III.2 Discrete Summation
We find it conceptually useful to begin by approximating an integral representation of the collective state by a discrete sum over $\omega_{i\mu}$, which approximates the continuous integrals in a large-$N$ limit. Define a collective field by summing over a large set of $N$ arbitrary $\omega_{i\mu}$ unit 4-vectors that span the future timelike hyperboloid:
$\Phi(r^{\mu})=\frac{1}{\sqrt{N}}\sum^{N}_{i=1}\phi_{\omega_{i}}(r^{\mu})\qquad\text{where,}\qquad\phi_{\omega_{i}}(r^{\mu})=\phi(\omega_{i}^{\mu}(\omega_{i\nu}r^{\nu})-r^{\mu})\qquad\omega_{i}^{2}=1.$ (25)
@@ III.2 Discrete Summation
$\widetilde{\Phi}(r^{\mu})=\frac{1}{\sqrt{N}}\Phi(r^{\mu})=\frac{1}{{N}}\sum^{N}_{i=1}\phi_{\omega_{i}}(r^{\mu});\qquad\widetilde{\Phi}(r^{\mu})=\phi_{\omega_{i}}(0)$ (26)
Consider the action for the single two-body bound state $H(X)\phi(r)$ of eqs.(75 to 79), replacing $\phi(r)\rightarrow\Phi(r)$ and $H(x)\rightarrow H^{\prime}(X)$. and denote renormalized parameters by primes . For the sake of discussion we break the action into separate components:
$S=S_{1}+S_{2}+S_{3}+S_{Y}+S_{\lambda}\qquad\text{where,}\qquad H(X)\phi(r)\rightarrow H^{\prime}(X)\Phi(r)$ (27)
@@
$S_{1}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(Z^{\prime}|DH^{\prime}(X)|^{2}|\Phi(r)|^{2}\biggr)$ (28)
$S_{2}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(Z^{\prime}|H^{\prime}(X)|^{2}|\partial_{r}\Phi(r)|^{2}\biggr)$ (29)
$S_{3}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(2{g^{\prime}{}_{0}^{2}N_{c}}\frac{1}{16M_{0}^{2}}\delta^{4}(r^{\mu})|H^{\prime\dagger}H^{\prime}||\Phi(r))|^{2}\biggr)$ (30)
@@
$D_{F}(2r^{\mu})\rightarrow\frac{1}{M_{0}^{2}}\delta^{4}(2r^{\mu})=\frac{1}{16M_{0}^{2}}\delta^{4}(r^{\mu}),$ (31)
@@
$\!\!\!\!\!\!\!\!\!\!S_{Y}=\widehat{g}^{\prime}_{Y}M_{0}^{2}\int\!\!d^{4}Xd^{4}r\;[\overline{\psi}_{iL}(X\!+\!r)\psi_{R}(X\!-\!r)]_{f}D_{F}(2r)\;H^{\prime i}(X)\Phi(\vec{r}){+h.c.}.\qquad\widehat{g}^{\prime}_{Y}\approx g^{\prime}{}_{0}^{2}\sqrt{2N_{c}/J}$ (32)
and the quartic interaction is given in the point-like limit of $\Phi(r)$:
@@
Our problem is to verify that $S(H^{\prime}\Phi)$, with renormalized parameters, is consistent with the underlying theory of the $\phi_{\omega}=\phi(r)$ as defined in eqs.(75,81). We therefore substitute the collective field definition of eq.25 and define the renormalized parameters.
@@
$\int\!\!d^{4}r\;\phi_{\omega_{i}}^{\dagger}(r)\phi_{\omega_{j}}(r)=\delta_{ij}\int\!\!d^{4}r\;\phi_{\omega}^{\dagger}(r)\phi_{\omega}(r)=\delta_{ij}$
$\int\!\!d^{4}r\;\partial_{\mu}\phi_{\omega_{i}}^{\dagger}(r)\partial^{\mu}\phi_{\omega_{j}}(r)=\delta_{ij}\int\!\!d^{4}r\;\partial_{\mu}\phi_{\omega}^{\dagger}(r)\partial^{\mu}\phi_{\omega}(r)$ (34)
@@
$S_{1}$: The new composite bilocal field takes the form $H^{\prime}(X)\Phi(r)$. The $H^{\prime}(X)$ kinetic term should be canonical with the Lorentz invariant $\Phi(r)$, hence we renormalize the $H^{\prime}(X)$ field as:
@@
This choice preserves the product $Z^{\prime}H^{\prime\dagger}H^{\prime}=ZH^{\dagger}H$, and $Z$ is still defined by the line integral relation of eq.(13) for the underlying $\phi_{\omega}$ field. We have from eqs.(25,28):
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S_{1}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(Z^{\prime}|DH^{\prime}(X)|^{2}|\Phi(r)|^{2}\biggr)=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r|DH^{\prime}|^{2}\biggl(\frac{Z^{\prime}}{N}\sum^{N}_{i=1}\sum^{N}_{j=1}\phi_{\omega_{i}}^{\dagger}(r)\phi_{\omega_{j}}(r)\biggr)$
@@
$ZM_{0}^{4}\int\!\!d^{4}r|\widetilde{\Phi}(r^{\mu})|^{2}=ZM_{0}^{4}\int\!\!d^{4}r|\phi_{\omega}(r^{\mu})|^{2}\;(\text{for any }\omega)=1$ (41)
$S_{1}$ is therefore consistent with the underlying two-body theory. Note that $\widetilde{\Phi}_{\omega}$ and $\phi_{\omega}$ integrates out with the normalization eq.(41).
@@
$\!\!\!\!\!\!\!\!\!\!S_{2}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(Z^{\prime}|H^{\prime}(X)|^{2}|\partial_{r}\Phi(r)|^{2}\biggr)=ZM_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(|H|^{2}\partial_{\mu}\phi_{\omega}^{\dagger}(r)\partial^{\mu}\phi_{\omega}(r)\biggr)\;(\text{for any }\omega)$
$=ZM_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(|H|^{2}\partial_{\mu}\widetilde{\Phi}^{\dagger}(r)\partial^{\mu}\widetilde{\Phi}(r)\biggr)$ (42)

### v3

@@ II Implementing the Relative Time Symmetry
We begin with the notion that the internal wave-function of two-body bound state, $\phi(r^{\mu})$, must not depend upon “relative time,” i.e., is independent of $r^{0}$ in the barycentric frame, or correspondingly $r^{\prime 0}$ in any frame. This can be seen in the free field limit by kinematics if we consider a pair of equal-mass particles of 4-momenta $p_{1}$ and $p_{2}$, $p_{1}^{2}=p_{2}^{2}=m^{2}$, and a bilocal wave-function, $H(x,y)\sim\exp(ip_{1}x+ip_{2}y)$. We pass to the total momentum $P=(p_{1}+p_{2})$ and relative momentum $Q=(p_{1}-p_{2})$, and the plane waves become $\exp(iPX+iQr)$, where we define “barycentric coordinates,” $x^{\mu}=X^{\mu}+r^{\mu}$, $y^{\mu}=X^{\mu}-r^{\mu}$. Note that $P_{\mu}Q^{\mu}=p_{1}^{2}-p_{2}^{2}=0$. Therefore, in the center-of-mass frame, in which $P=(P^{0},\vec{0})$ and $Q=(0,\vec{q})$, we see that $Q^{0}=0$. This implies there is no dependence in the bilocal state on $r^{0}$ through $\exp(iQ_{0}r^{0})=1$, and likewise no dependence upon a boosted relative time $r^{\prime 0}$ in any other frame. If the particles are constituents of a bound state, then the “relative time” must decouple from the dynamics. Indeed, it is useful to think of the bound state as free field pair of particles for which the bilocal field is properly normalized, and the interaction is subsequently adiabatically switched on.
@@
Given an arbitrary Lorentz invariant function, $\phi(r^{\mu})$, in any frame there will generally be dependence upon a relative time, $\sim\tau$. This is analogous to the gauge dependent components of a vector potential, and it is an artifact of using the bilocal field description. [^3] I am grateful to Bill Bardeen for some discussions that inspired this perspective. As described in the Introduction, the relative time, $\tau$, can be written in any given frame as $r^{\mu}=\omega^{\mu}\tau$, and implicitly requires the timelike, 4-vector, $\omega^{\mu}$, the “arrow of relative time.” Hence, eliminating dependence upon $\tau$, requires a Lorentz invariant constraint, such as $\omega_{\mu}\partial^{\mu}\phi(r)=0$. In the symmetric phase of the standard model (SM), (or for any typical two-body bound state) the BEH boson contains such a vector, i.e., the 4-momentum $P^{\mu}$ carried by $H(X^{\mu})\sim\exp(iP_{\mu}X^{\mu})$. We can therefore bootstrap $\omega^{\mu}$ to $P^{\mu}$ through a constraint relation $\omega^{\mu}\propto P^{\mu}$. For example, we can do this semiclassically by introducing Lagrange multipliers into the action, such as:
$W=M_{0}^{4}\int d^{4}Xd^{4}r\;\lambda^{\prime}(H^{\dagger}D_{\mu}H\phi^{\dagger}\partial_{\mu}\phi),$ (8)
@@
where we demand $\delta W/\delta\lambda^{\prime}=0$, which imposes the kinematic constraint $P_{\mu}Q^{\mu}=0$. However, the question then remains: “what happens in the vacuum where $P_{\mu}=0$ and $\omega^{\mu}$ becomes unconstrained?”
@@
$\!\!\!\!\!\!\!\!\!\!S_{\phi}$ $\rightarrow$ $M_{0}^{4}\int\!\!d^{4}r\biggl(Z\partial_{\mu}\phi_{\omega}^{\dagger}(r)\partial^{\mu}\phi_{\omega}(r)+2g_{0}^{2}N_{c}D_{F}(2r^{\mu})\phi_{\omega}^{\dagger}(r)\phi_{\omega}(r)\biggr),$ (10)
@@
$M_{0}\int\!\!\biggl(-Z\frac{\partial^{2}\phi_{\omega}(r^{\mu})}{\partial r^{\mu}\partial r_{\mu}}+2g_{0}^{2}N_{c}D_{F}(2r^{\mu})\phi_{\omega}(r^{\mu})\biggr)\omega_{\nu}dr^{\nu}=ZM_{0}\int\!\!\mu^{2}\phi_{\omega}(r^{\mu})\omega_{\nu}dr^{\nu}=\mu^{2}\phi_{\omega}(r^{\mu}).$ (11)
Note the presence of the overall line integral, $\int\omega_{\mu}dr^{\mu}$. This line integral remains in the equation of motion since $\phi_{\omega}(r^{\mu})$ has no dependence upon $r_{\mu}\propto\omega_{\mu}$, hence the variation is constrained, $\delta\phi_{\omega}\sim\delta^{3}(r^{\mu}_{\perp})$ where $\omega_{\mu}r^{\mu}_{\perp}=0$ and does not produce a longitudinal variation, $\delta(\omega_{\mu}r^{\mu}$).
We define $Z$ by the line integral normalization chris1 ,
@@
$Z$ is defined to canonically normalize the free field pair of particles before the interaction is turned on. It is important to realize that eq.(11) is not a conventional Klein-Gordon equation due to the line integral constraint.
@@
$\!\!\!\!\!\!\!\!\!\!S_{\phi}=M_{0}^{3}\!\!\int\!\!d^{3}r\;\biggl(-|\nabla_{\vec{r}}\phi(\vec{r})|^{2}+\int dr^{0}\;{2g_{0}^{2}N_{c}M_{0}}D_{F}(2r^{\mu})|\phi(\vec{r})\;|^{2}\biggr).$ (13)
@@
In the solution to the SKG equations with the eigenvalue, $\mu^{2}$, we see upon integrating by parts that $S_{\phi}\rightarrow\mu^{2}$, and the full bound state action then becomes (in the pointlike limit of the interaction):
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S=\!\!\int\!\!d^{4}X\;\biggl(|D_{H}H(X^{\mu})|^{2}-\mu^{2}|H(X^{\mu})|^{2}-\frac{\lambda}{2}(H^{\dagger}H)^{2}-g_{Y}\left([\overline{\psi}_{iL}(X)t_{R}(X)]_{f}H^{i}(X)+h.c.\right)\biggr)+S^{\prime}.$ (17)
$\mu^{2}$ is the physical mass of the bound state. The Yukawa coupling, $g_{Y}$ and quartic coupling $\lambda$ are defined in eqs.(85-90) in Appendix B. We emphasize that, while we have evaluated the action in the particular $\omega^{\mu}$ frame in eq.(13), this is just a calculation in a simplifying frame in the overall Lorentz invariant action eq.(11). The result eq.(17) holds in any frame.
@@
$\frac{g_{c}^{2}N_{c}}{8\pi^{2}}=1.06940,\qquad\makebox{c.f, the NJL critical value, }\qquad\left.\frac{g_{0}^{2}N_{c}}{8\pi^{2}}\right|_{NJL}=1.00.$ (18)
@@
$\overline{g}_{0}^{2}=g_{0}^{2}\left(1-\frac{g_{0}^{2}N_{c}}{8\pi^{2}}\right)^{-1}.$ (19)
When $\overline{g}_{0}>g_{c}$ the eigenvalue is $\mu^{2}<0$. In such a solution the action eq.(17) for $H(X^{\mu})$ then yields the “sombrero potential”:
@@ III Broken Phase, Manifestly Lorentz Invariant Vacuum, and BEH Excitations
Presently we show how the action can be written in terms of the collective field $\Phi(r^{\mu})$. We find it conceptually useful to begin by approximating an integral representation of the collective state by a discrete sum over $N$ of the solutions, $\omega_{i\mu}$. This approximates the continuous integrals which are defined subsequently corresponding to the large-$N$ limit. We begin with a brief discussion of the orthogonality of the $\phi_{\omega_{i}}$ which is further treated in Appendix A. In Appendix B for reference, we give a formal summary of the natural top condensation scheme.
@@ III.1 Formal Derivation
The action of eq.(10) is manifestly Lorentz invariant with dependence upon the 4-vector $\omega^{\mu}$. For the spherically symmetric ground state it must yield a Lorentz invariant expression in $\omega$. Hence, under a Lorentz transformation, $\omega^{\prime}_{\mu}=\Lambda_{\mu}^{\nu}\omega^{\nu}$ we can likewise perform $r^{\prime}_{\mu}=\Lambda_{\mu}^{\nu}r^{\nu}$, however, we are then free to change the integration variable back to its original form, $r^{\prime}\rightarrow r$ to obtain:
$\!\!\!\!\!\!\!\!\!\!S_{\phi}=M_{0}^{4}\int\!\!d^{4}r\biggl(Z\partial_{\mu}\phi_{\omega^{\prime}}^{\dagger}(r)\partial^{\mu}\phi_{\omega^{\prime}}(r)+2g_{0}^{2}N_{c}D_{F}(2r^{\mu})\phi_{\omega^{\prime}}^{\dagger}(r)\phi_{\omega^{\prime}}(r)\biggr).$ (21)
@@
$ZM_{0}^{4}\int\!\!d^{4}r\;\phi_{\omega}^{\dagger}(r^{\mu})F(r^{\mu})\phi_{\omega`}(r^{\mu})=ZM_{0}^{4}\int\!\!d^{4}r\;\delta^{4}(\omega-\omega^{\prime})\phi_{\omega}^{\dagger}(r^{\mu})F(r^{\mu})\phi_{\omega`}(r^{\mu}).$ (22)
@@
$ZM_{0}^{4}\int\!\!d^{4}r\;\phi_{\omega}^{\dagger}(r^{\mu})\delta^{4}(r^{\mu})\phi_{\omega`}(r^{\mu})=ZM_{0}^{4}\phi_{\omega}^{\dagger}(0)\phi_{\omega`}(0)=ZM_{0}^{4}\int\!\!d^{4}r\;\delta^{4}(r^{\mu})\phi_{\omega}^{\dagger}(r)\phi_{\omega`}(r).$ (23)
This implies that we will have the BCS or BEC-like behavior for $\Phi(r^{\mu})$. We derive the orthogonality in Appendix A.
Define a collective field by summing over a large set of $N$ arbitrary $\omega_{i\mu}$ unit 4-vectors that span the future timelike hyperboloid:
$\Phi(r^{\mu})=\frac{1}{\sqrt{N}}\sum^{N}_{i=1}\phi_{\omega_{i}}(r^{\mu}),\qquad\text{where,}\qquad\phi_{\omega_{i}}(r^{\mu})=\phi(\omega_{i}^{\mu}\omega_{i}^{\nu}r_{\nu}-r^{\mu}),\qquad\omega_{i}^{2}=1.$ (24)
@@
$\widetilde{\Phi}(r^{\mu})=\frac{1}{\sqrt{N}}\Phi(r^{\mu})=\frac{1}{{N}}\sum^{N}_{i=1}\phi_{\omega_{i}}(r^{\mu});\qquad\widetilde{\Phi}(r^{\mu})=\phi_{\omega_{i}}(0).$ (25)
Consider the action for the single two-body bound state $H(X^{\mu})\phi(r)$ of eqs.(81 to 85), replacing $\phi(r)\rightarrow\Phi(r^{\mu})$ and $H(x)\rightarrow H^{\prime}(X)$. and denote renormalized parameters by primes . For the sake of discussion we break the action into separate components:
$S=S_{1}+S_{2}+S_{3}+S_{Y}+S_{\lambda},\qquad\text{where,}\qquad H(X^{\mu})\phi(r)\rightarrow H^{\prime}(X)\Phi(r^{\mu}),$ (26)
@@
$S_{1}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(Z^{\prime}|DH^{\prime}(X)|^{2}|\Phi(r^{\mu})|^{2}\biggr),$ (27)
$S_{2}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(Z^{\prime}|H^{\prime}(X)|^{2}|\partial_{r}\Phi(r^{\mu})|^{2}\biggr),$ (28)
$S_{3}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(2{g^{\prime}{}_{0}^{2}N_{c}}\frac{1}{16M_{0}^{2}}\delta^{4}(r^{\mu})|H^{\prime\dagger}H^{\prime}||\Phi(r^{\mu}))|^{2}\biggr),$ (29)
@@
$D_{F}(2r^{\mu})\rightarrow\frac{1}{M_{0}^{2}}\delta^{4}(2r^{\mu})=\frac{1}{16M_{0}^{2}}\delta^{4}(r^{\mu}).$ (30)
@@
$\!\!\!\!\!\!\!\!\!\!S_{Y}=\widehat{g}^{\prime}_{Y}M_{0}^{2}\int\!\!d^{4}Xd^{4}r\;[\overline{\psi}_{iL}(X\!+\!r)\psi_{R}(X\!-\!r)]_{f}D_{F}(2r)\;H^{\prime i}(X)\Phi(\vec{r}){+h.c.},\qquad\widehat{g}^{\prime}_{Y}\approx g^{\prime}{}_{0}^{2}\sqrt{2N_{c}/J},$ (31)
and the quartic interaction is given in the point-like limit of $\Phi(r^{\mu})$:
@@
Our problem is to verify that $S(H^{\prime}\Phi)$, with renormalized parameters, is consistent with the underlying theory of the $\phi_{\omega}=\phi(r)$ as defined in eqs.(81,87). We therefore substitute the collective field definition of eq.24 and define the renormalized parameters.
@@
$\int\!\!d^{4}r\;\phi_{\omega_{i}}^{\dagger}(r)\phi_{\omega_{j}}(r)=\delta_{ij}\int\!\!d^{4}r\;\phi_{\omega}^{\dagger}(r)\phi_{\omega}(r),$
$\int\!\!d^{4}r\;\partial_{\mu}\phi_{\omega_{i}}^{\dagger}(r)\partial^{\mu}\phi_{\omega_{j}}(r)=\delta_{ij}\int\!\!d^{4}r\;\partial_{\mu}\phi_{\omega}^{\dagger}(r)\partial^{\mu}\phi_{\omega}(r).$ (33)
@@
Note that these normalizations differ from the BCS superconductor in two major regards. First, $Z$ is treated as an extensive parameter, i.e., when we renormalize each $\phi_{\omega}\rightarrow(1/\sqrt{N})\phi_{\omega}$ to comprise the sum over $\omega^{\mu}$ we then redefine $Z\rightarrow Z^{\prime}=ZN$. The new composite bilocal field then takes the form $H^{\prime}(X)\Phi(r^{\mu})$.
Secondly, the $H^{\prime}(X)$ kinetic term should be canonical with the Lorentz invariant $\Phi(r^{\mu})$, hence we renormalize the $H^{\prime}(X)$ field as:
@@
This choice preserves the product $Z^{\prime}H^{\prime\dagger}H^{\prime}=ZH^{\dagger}H$, and $Z$ is still defined by the line integral relation of eq.(12) for the underlying $\phi_{\omega}$ field. These normalizations will yield the same predictions for $g^{2}_{critical}$ and $M_{0}\sim 6$ TeV, as one would obtain for a single $\phi_{\omega}$ condensate.
We now discuss the calculation of the full action term by term:
$S_{1}$: We have from eqs.(24,27):
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S_{1}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(Z^{\prime}|DH^{\prime}(X)|^{2}|\Phi(r^{\mu})|^{2}\biggr)=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r|DH^{\prime}|^{2}\biggl(\frac{Z^{\prime}}{N}\sum^{N}_{i=1}\sum^{N}_{j=1}\phi_{\omega_{i}}^{\dagger}(r)\phi_{\omega_{j}}(r)\biggr)$
@@
$ZM_{0}^{4}\int\!\!d^{4}r|\widetilde{\Phi}(r^{\mu})|^{2}=ZM_{0}^{4}\int\!\!d^{4}r|\phi_{\omega}(r^{\mu})|^{2}\;(\text{for any }\omega)=1.$ (40)
$S_{1}$ is therefore consistent with the underlying two-body theory. Note that $\widetilde{\Phi}_{\omega}$ and $\phi_{\omega}$ integrate out with the normalization eq.(40).
@@
$\!\!\!\!\!\!\!\!\!\!S_{2}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(Z^{\prime}|H^{\prime}(X)|^{2}|\partial_{r}\Phi(r^{\mu})|^{2}\biggr)=ZM_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(|H|^{2}\partial_{\mu}\phi_{\omega}^{\dagger}(r)\partial^{\mu}\phi_{\omega}(r)\biggr)\;(\text{for any }\omega)$
$=ZM_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(|H|^{2}\partial_{\mu}\widetilde{\Phi}^{\dagger}(r)\partial^{\mu}\widetilde{\Phi}(r)\biggr),$ (41)